A student has carried out a moving boundary experiment using K2SO4 as an electrolyte.

The data obtained was tabulated in the following Table.

Cation velocity 11.43 mms-1

Cross sectional area of tube 28.56 mm2

Transference number of cation 0.593

Anion concentration 0.025 mol dm-3

Calculate the current flow for the boundary moves as far as 45 mm


1
Expert's answer
2021-01-11T03:49:24-0500

The transference number of cation (t+) is defined as the ratio of the electric current derived from the cation to the total electric current.

t+ = "\frac{z_+*c*L*A*F}{Q}"

"Q=\frac{z_+*c*L*A*F}{t_+}"

z+ = 1 (charge of cation)

c(K+) = 0.025*2 = 0.05 mol/dm3 = 0.00005 mol/cm3

L = 45 mm = 4.5 cm

A (cross-sectional area of tube) = 28.56 mm2 = 0.2856 cm2

F = 96500 C/mol

t+ = 0.593

"Q=\frac{1*0.00005*4.5*0.2856*96500}{0.593}=10.46" C